# MLPs Learn In-Context

### v1

@@ A.2 A different input scheme
Is there a way to circumvent this impossibility result? One aspect of the proof that may seem unconventional is its reliance on flattening one-hot encodings $\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}))$ as input to the MLP. Going as far back as Word2vec [38], the established convention for processing one-hot inputs is to instead pass them through an embedding matrix $\boldsymbol{W}_{e}$, creating vector embeddings
@@
where $d$ is the dimension of a single symbol’s vector embedding. A practitioner then operates on the resulting vector embeddings, not the one-hot encodings directly. As we will shortly see, if we consider the more conventional input scheme that uses vector embeddings $\boldsymbol{h}_{0}(\boldsymbol{x})$ and not the one-hot encodings directly, then the conclusion from Boix-Adsera et al. [33] no longer holds.
@@
In this way, Theorem A.1 can still be rescued through permutations on one-hots before the embedding matrix. This method sidesteps the issue with permuting $\mathsf{flat}(\boldsymbol{h}_{0}(\boldsymbol{x}))$ directly, and the MLP trained on SGD remains invariant to any permutation on the underlying one-hots. Hence, it seems the proof remains valid, and the impossibility result still holds. As an aside, permuting over individual one-hot vectors is equivalent to permuting over the underlying symbols. Hence, for notational clarity, we will redefine $\Pi$ to be a permutation over symbols $\Pi:\mathcal{X}\rightarrow\mathcal{X}$, and write $\Pi\boldsymbol{x}=\Pi(x_{1})\Pi(x_{2})\ldots\Pi(x_{k})$ by bijectively mapping each symbol in $\mathcal{X}$ to a different symbol in $\mathcal{X}$. This allows us to write the vector embeddings over permuted one-hots as
$\boldsymbol{h}_{0}^{\pi}(\boldsymbol{x})=\boldsymbol{h}_{0}(\Pi\boldsymbol{x})\,.$
Unfortunately, this scheme runs into a different issue: it is impossible to find two inputs $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}$ where $\Pi\boldsymbol{x}^{(1)}=\boldsymbol{x}^{(2)}$, but that $f^{*}(\boldsymbol{x}^{(1)})\neq f^{*}(\boldsymbol{x}^{(2)})$. Indeed, we next show that if $\boldsymbol{x}$ satisfies a template $\boldsymbol{z}$, then any permutation $\Pi$ on the symbols of $\boldsymbol{x}$ will also satisfy $\boldsymbol{z}$. This can be seen quite simply by considering that 1) by definition, template satisfaction is invariant under a relabeling of symbols and 2) any permutation is a relabeling of symbols — hence, template satisfaction must be invariant under permutation. We phrase this formally below.
@@ A.3 Can MLPs reason relationally?
In this way, if a model satisfies the above three conditions, it generalizes to unseen inputs for a task defined by the labeling function $f^{*}$. The first condition for smoothness is regularly achieved by standard neural networks [39]. The second condition corresponds to a notion of expressivity — that is, a model $f$ should be able to approach arbitrarily close to zero on its training data. For modern neural network models trained on simple tasks, this is a frequent occurrence [40]. The third condition corresponds to a coarse description of data diversity. The training data should be sufficiently diverse such that all unseen examples are very close to an example seen during training. This condition may be difficult to achieve in practice, but it offers a very coarse upper bound on the requisite data diversity required to generalize on unseen examples. Nonetheless, an MLP trained online on a suitably constrained input space may very well achieve this condition with sufficiently many training iterations. Given further assumptions on $f$ (e.g. $f$ is an MLP), it is likely possible to shrink this data diversity bound considerably.
@@ A.5 Discussion
Despite our success above, many earlier studies document cases where common neural network architectures fail to reason relationally [41, 42, 43, 32]. One important reason for the failure may be that the task inputs are very large and complex, as in visual reasoning [42, 44]. Proposition A.2 suggests that the data diversity required for successful generalization scales exponentially with the dimension of the inputs in the worst case. It is possible that given a sufficiently vast dataset, an MLP would perform well on visual reasoning tasks. Furthermore, having shown above that MLPs are decisively capable of relational reasoning (especially when presented with idealized stimulus embeddings, as in Section 3), their failure on complex tasks highlights a need to separate a model’s ability to reason relationally from its ability to learn sufficiently rich feature representations. In realistic data-limited scenarios, perhaps an MLP paired with a more bespoke module for feature learning would reason quite successfully. We anticipate further work that more closely investigates whether these failures stem from data limitations, insufficient feature learning, or some other cause, thereby building a more complete and updated picture of relational reasoning in neural networks.
@@ B.3 Discussion
Evidently simple tasks with long inputs work against the Transformer’s attention mechanism. Shortening the context by reducing the task dimension, chunking inputs into larger tokens, or bypassing the attention mechanism altogether by stacking the input into a single token all improve the Transformer’s efficiency. It is not immediately obvious why the Transformer performs so dramatically worse compared to the MLP for larger $n$, though it is well-known that Transformers can struggle with long inputs [45]. These experiments add one more example supporting the bitter lesson [13] that less inductive bias is better.
@@ C.2 MLP
For all tasks, we use ReLU activation functions applied pointwise $\phi(\boldsymbol{x})=\max(\boldsymbol{x},\boldsymbol{0})$. Widths of all hidden layers are fixed to the same value $H$. As with all models, all training examples are presented online with batch size 128. Training uses AdamW [46] with learning rate $\alpha=1\times 10^{-4}$ and weight decay $\lambda=1\times 10^{-4}$. The hyperparameters used to train MLPs on each task are presented in Table 1.
@@ C.6 ICL regression
We prepare in-context regression in a setup that closely mimics Raventós et al. [15], though without an autoregressive objective. The input consists of a sequence of values $(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\ldots,(\boldsymbol{x}_{L},y_{L})$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$. The $\boldsymbol{x}_{i},y_{i}$ pairs are linearly related through a set of weights $\boldsymbol{\beta}\in\mathbb{R}^{n}$ such that $y_{i}=\boldsymbol{x}_{i}\cdot\boldsymbol{\beta}+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ corresponds to noise. Finally, the input includes a query $\boldsymbol{x}_{q}$. The model output is a single scalar regressed against the corresponding $y_{q}$. Inputs are sampled as $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and weights are sampled as $\boldsymbol{\beta}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}/n\right)$. Before being presented to the model, all inputs are packed into an input matrix $\tilde{\boldsymbol{X}}\in\mathbb{R}^{(L+1)\times(n+1)}$ with the following structure [8]
@@ C.14 Software
All models are implemented and trained using the Jax [48] family of libraries, particularly Flax [49]. Plots are created using Seaborn [50] and Pandas [51].
@@ Appendix D Additional figures
Figure 6: ICL classification supplementary figures. (a) MSE obtained across each architecture as a function of compute. Lines connect common models, with colors denoting different parameter counts. Hence, a single line traces the trajectory of a model across different training iterations. (b) Cross entropy loss across different context lengths $L$, for different input dimensions $n$. Line colors indicate the number of elapsed training steps. In these examples, $B=n/2$, so there are only 2 labels present in each context (out of $C=32$ total possible labels). The gray dashed line corresponds to the loss obtained by placing equal probability on the 2 labels present in the context. All models plateau for a time at guessing one among the two correct labels, before eventually collapsing to the correct ICL solution. (c) IWL to ICL transition for different burstiness $B$. Consistent with prior work [9, 16], higher burstiness encourages ICL. Transformers transition to ICL for lower burstiness and lower number of clusters $k$. (d) ICL vs. IWL behavior for $B=n/2$ and $k=2048$ clusters across context lengths $L$ and input dimensions $n$. For the most part, these settings are sufficient to encourage ICL, including the configuration plotted in the main text Figure 1, though ICL appears to decay at higher dimensions and longer contexts. (all) Line plots feature 95 percent confidence intervals about the mean, computed across 5 replications.

### v3

@@ A.2 A different input scheme
Is there a way to circumvent this impossibility result? One aspect of the proof that may seem suspect is its reliance on flattening one-hot encodings $\mathsf{flat}(\boldsymbol{E}(\boldsymbol{x}))$ as direct input to the MLP. Going as far back as Word2vec [Mikolov et al., 2013], a well-established convention for processing one-hot inputs is to instead pass them through an embedding matrix $\boldsymbol{W}_{e}$, creating vector embeddings
@@
where $d$ is the dimension of a single symbol’s vector embedding. [^5] Indeed, in their results on Transformers, Boix-Adsera et al. do use vector embeddings. It is unusual they would choose to omit them in their analysis of MLPs. A practitioner then flattens and operates on the resulting vector embeddings, not the one-hot encodings directly. As we will shortly see, if we consider the more conventional input scheme that uses vector embeddings $\boldsymbol{h}_{0}(\boldsymbol{x})$ and not the one-hot encodings directly, then the conclusion from Boix-Adsera et al. [2023] no longer holds.
@@
In this way, Theorem A.1 might still be rescued through permutations on one-hots before the embedding matrix. This method sidesteps the issue with permuting $\mathsf{flat}(\boldsymbol{h}_{0}(\boldsymbol{x}))$ directly, and the MLP trained on SGD remains invariant to any permutation on the underlying one-hots. Hence, it seems the proof may remain valid, and the impossibility result might still holds.
Unfortunately, this scheme runs into a different issue: it is impossible to find two inputs $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)}$ where $\Pi\boldsymbol{x}^{(1)}=\boldsymbol{x}^{(2)}$, but that $f^{*}(\boldsymbol{x}^{(1)})\neq f^{*}(\boldsymbol{x}^{(2)})$. (Note, we have abused notation slightly and write $\Pi\boldsymbol{e}_{x}=\Pi\boldsymbol{x}$.) Indeed, we next show that if $\boldsymbol{x}$ satisfies a template $\boldsymbol{z}$, then any permutation $\Pi$ on the symbols of $\boldsymbol{x}$ will also satisfy $\boldsymbol{z}$. This can be seen quite simply by considering that 1) by definition, template satisfaction is invariant under a relabeling of symbols and 2) any permutation is a relabeling of symbols — hence, template satisfaction must be invariant under permutation. We phrase this formally below.
@@ A.3 Can MLPs reason relationally?
In this way, if a model satisfies the above three conditions, it generalizes to unseen inputs for a task defined by the labeling function $f^{*}$. The first condition for smoothness is regularly achieved by standard neural networks [Khromov and Singh, 2023]. The second condition corresponds to a notion of expressivity — that is, a model $f$ should be able to approach arbitrarily close to zero on its training data. For modern neural network models trained on simple tasks, this is a frequent occurrence [Zhang et al., 2021]. The third condition corresponds to a coarse description of data diversity. The training data should be sufficiently diverse such that all unseen examples are very close to an example seen during training. This condition may be difficult to achieve in practice, but it offers a very coarse upper bound on the requisite data diversity required to generalize on unseen examples. Nonetheless, an MLP trained online on a suitably constrained input space may very well achieve this condition. Given further assumptions on $f$ (e.g. $f$ is an MLP), it is likely possible to shrink this data diversity bound considerably.
@@ A.5 Discussion
Despite our success above, many earlier studies document cases where common neural network architectures fail to reason relationally [Marcus et al., 1999, Kim et al., 2018, Lake and Baroni, 2018, Alhama and Zuidema, 2019]. One important reason for the failure may be that the task inputs are very large and complex, as in visual reasoning [Kim et al., 2018, Serre, 2019]. Proposition A.2 suggests that the data diversity required for successful generalization scales exponentially with the dimension of the inputs in the worst case. It is possible that given a sufficiently vast dataset, an MLP would perform well on visual reasoning tasks. Furthermore, having shown above that MLPs are decisively capable of relational reasoning (especially when presented with idealized stimulus embeddings, as in Section 3), their failure on complex tasks highlights a need to separate a model’s ability to reason relationally from its ability to learn sufficiently rich feature representations. In realistic data-limited scenarios, perhaps an MLP paired with a more bespoke module for feature learning would reason quite successfully. We anticipate further work that more closely investigates whether these failures stem from data limitations, insufficient feature learning, or some other cause, thereby building a more complete and updated picture of relational reasoning in neural networks.
@@ B.3 Discussion
Evidently simple tasks with long inputs work against the Transformer’s attention mechanism. Shortening the context by reducing the task dimension, chunking inputs into larger tokens, or bypassing the attention mechanism altogether by stacking the input into a single token all improve the Transformer’s efficiency. It is not immediately obvious why the Transformer performs so dramatically worse compared to the MLP for larger $n$, though it is well-known that Transformers can struggle with long inputs [Tay et al., 2020].
@@ C.2 MLP
For all tasks, we use ReLU activation functions applied pointwise $\phi(\boldsymbol{x})=\max(\boldsymbol{x},\boldsymbol{0})$. Widths of all hidden layers are fixed to the same value $H$. As with all models, all training examples are presented online with batch size 128. Training uses AdamW [Loshchilov and Hutter, 2017] with learning rate $\alpha=1\times 10^{-4}$ and weight decay $\lambda=1\times 10^{-4}$. The hyperparameters used to train MLPs on each task are presented in Table 1.
@@ C.6 ICL regression
We prepare in-context regression in a setup that closely mimics Raventós et al. [2024], though without an autoregressive objective. The input consists of a sequence of values $(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\ldots,(\boldsymbol{x}_{L},y_{L})$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$. The $\boldsymbol{x}_{i},y_{i}$ pairs are linearly related through a set of weights $\boldsymbol{\beta}\in\mathbb{R}^{n}$ such that $y_{i}=\boldsymbol{x}_{i}\cdot\boldsymbol{\beta}+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ corresponds to noise. Finally, the input includes a query $\boldsymbol{x}_{q}$. The model output is a single scalar regressed against the corresponding $y_{q}$. Inputs are sampled as $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and weights are sampled as $\boldsymbol{\beta}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}/n\right)$. Before being presented to the model, all inputs are packed into an input matrix $\tilde{\boldsymbol{X}}\in\mathbb{R}^{(L+1)\times(n+1)}$ with the following structure [Zhang et al., 2023]
@@ C.14 Software
All models are implemented and trained using the Jax [Bradbury et al., 2018] family of libraries, particularly Flax [Heek et al., 2023]. Plots are created using Seaborn [Waskom, 2021] and Pandas [pandas development team, 2020].
@@ Appendix D Additional figures
Figure 5: ICL regression with an autoregressive objective. For each input example $(\boldsymbol{x}_{1},y_{1},\boldsymbol{x}_{2},y_{2},\ldots,\boldsymbol{x}_{L},y_{L})$, we compute the autoregressive loss $\sum_{i}\mathcal{L}(f(\boldsymbol{x}_{1},y_{1},\boldsymbol{x}_{2},y_{2},\ldots\boldsymbol{x}_{i}),y_{i})$, for a neural network $f$ and MSE loss $\mathcal{L}$. For vanilla MLPs and Mixers, variable-length inputs are handled by padding inputs with zero up to the max length $L$. (a) Compute vs. MSE on the unrestricted task distribution. Each point represents a single model, with particular parameters and training iterations. Just as in the fixed input length case, at large compute, MSE is approximately equal across all architectures. The red line corresponds to the Bayes optimal Ridge MSE. (b) Excess MSE (MSE above Bayes optimal) for varying context length $L$ on the unrestricted task distribution. Excess MSE remains flat for Mixers and Transformers, but rises for MLPs. The grey line corresponds to the excess MSE incurred by the zero predictor. Given compute limitations, we plot on a slightly narrower range of context lengths, but the overall trends remain consistent with the finite-input-length case. (c, d) IWL to ICL transition with increasing data diversity. We train on a finite distribution with $k$ weights, then test on both the finite training distribution and the unrestricted distribution. Just as with finite input lengths, all models exhibit a transition from IWL (represented by dMMSE) to ICL (represented by Ridge) as $k$ increases. Note: it is possible to “outperform" Bayes optimal Ridge on the finite training distribution by learning in-weight the underlying $\boldsymbol{\beta}$’s. (all) We use $n=8$ dimension inputs. All line plots feature 95 percent confidence intervals about the mean, estimated from 5 replications.
@@
Figure 7: ICL classification supplementary figures. (a) MSE obtained across each architecture as a function of compute. Lines connect common models, with colors denoting different parameter counts. Hence, a single line traces the trajectory of a model across different training iterations. (b) Cross entropy loss across different context lengths $L$, for different input dimensions $n$. Line colors indicate the number of elapsed training steps. In these examples, $B=n/2$, so there are only 2 labels present in each context (out of $C=32$ total possible labels). The gray dashed line corresponds to the loss obtained by placing equal probability on the 2 labels present in the context. All models plateau for a time at guessing one among the two correct labels, before eventually collapsing to the correct ICL solution. (c) IWL to ICL transition for different burstiness $B$. Consistent with prior work [Reddy, 2024, Chan et al., 2022], higher burstiness encourages ICL. Transformers transition to ICL for lower burstiness and lower number of clusters $k$. (d) ICL vs. IWL behavior for $B=n/2$ and $k=2048$ clusters across context lengths $L$ and input dimensions $n$. For the most part, these settings are sufficient to encourage ICL, including the configuration plotted in the main text Figure 1, though ICL appears to decay at higher dimensions and longer contexts. (all) Line plots feature 95 percent confidence intervals about the mean, computed across 5 replications.